OpenStudy (anonymous):

solve for x. (2x + 60) + 3x = 180

OpenStudy (anonymous):

180 + 60 = 240

OpenStudy (anonymous):

umm.. i mean180-60=2x+3x

OpenStudy (anonymous):

120

OpenStudy (anonymous):

2x+3x= 5x

OpenStudy (anonymous):

um i still dont know how to get x

OpenStudy (anonymous):

2x+3x=180-60 5x=120 x=120/5 SOLVE IT :D